Use the following information on states of the economy and stock returns to calculate the standard deviation of returns. Assume that all three states are equally likely. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

State of Economy Security Return if State Occurs

Recession −6.50 %

Normal 9.00

Boom 19.00
Answer #1

Expected Return=Respective return*respective probability

=(1/3*-6.5)+(1/3*9)+(1/3*19)

=7.167%(Approx).

probability Return probability*(Return-Expected return)^2
1/3 -6.5 1/3*(-6.5-7.167)^2=62.25925926
1/3 9 1/3*(9-7.167)^2=1.12037037
1/3 19 1/3*(19-7.167)^2=46.67592593
Total=110.056%(Approx).

Standard deviation=[Total probability*(Return-Expected return)^2/Total probability]^(1/2)

=(110.056)^(1/2)

=10.49%(Approx).